OpenStudy (anonymous):

Find a polynomial of degree 4 that has integer coefficients and zeros 2, −2, 3, and 1/2 .

OpenStudy (anonymous):

HELP ASAP

OpenStudy (mathstudent55):

If a polynomial with integer coefficients has roots a, b, c, d, then the polynomial is (x - a)(x - b)(x - c)(x - d)

jimthompson5910 (jim_thompson5910):

x = a/b bx = a bx - a = 0 -a + bx = 0 a - bx = 0 The work above shows that if x = a/b is a root of the polynomial, then (a-bx) is a factor of the polynomial